In a certain city, the daily consumption of water (in millions of litres) is a random variable with probability density function given by:

𝟏 π’†βˆ’π’™ 𝒇(𝒙)={πŸ—πŸ— , 𝒙β‰₯𝟎

𝟎

(i) Identify the distribution and use its parameter to find the city’s expected water consumption for any given day. [3 marks]

(ii) What is the probability that the water supply on a given day is inadequate if the daily capacity of the city is 9 million litres?


1
Expert's answer
2021-06-14T09:13:28-0400

Let "m= \frac{1}{9}"

"f(x) = me^{-xm}, \;x≥0"

(i) This distribution is exponential with mean "= \frac{1}{m}=9"

Expected water consumption = 9 million liters.

(ii) CDF of exponential distribution "= P(X\u2264x)= 1 -e^{-mx}"

For inadequate supply, daily capacity should be less than 9 million liters.

"P(X≤9) = 1 -e^{- \frac{1}{9} \times 9} = 0.632"